OpenStudy (anonymous):

Find the median of the following data set. 1 1/4,5/8,3/5 ,1/2 , 1 1/2, 1 3/4

OpenStudy (anonymous):

answers 5/6 7/8 15/16

OpenStudy (mathteacher1729):

Do you know the definition of the word median?

OpenStudy (anonymous):

middle

OpenStudy (mathteacher1729):

Yes. So the key is to write all of these as fractions, with the same denominator.

OpenStudy (mathteacher1729):

Then just find the one that's in the middle.

OpenStudy (anonymous):

ok how do i do that

OpenStudy (mathteacher1729):

Let's start with the very first number: 1 1/4 that means 1 plus 1/4 What is that as a fraction?

OpenStudy (anonymous):

5/4

OpenStudy (mathteacher1729):

Good. :) The only other mixed number there is 1 1/2 (That's 1 plus 1/2)

OpenStudy (anonymous):

or 3/2

OpenStudy (anonymous):

and the last one which would be 7/4

OpenStudy (mathteacher1729):

Yes. So now you have a bunch of fractions. 5/4 ,5/8 , 3/5 , 1/2 , 3/2 , oh wait ... 1 3/4

OpenStudy (anonymous):

7/4

OpenStudy (mathteacher1729):

Yes, so now the set it this... 5/4 5/8 3/5 1/2 3/2 7/4 Check out those denominators... 2, 4, 8, 5 What is the lowest common multiple of 2,4,8, and 5?

OpenStudy (anonymous):

1

OpenStudy (mathteacher1729):

You found the greatest common factor... :)

OpenStudy (anonymous):

ok so whats next then

OpenStudy (mathteacher1729):

You have to give all the fractions a common denominator. That is done by finding the lowest common multiple. For example, how would you give 1/2 and 1/3 the same denominator?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i do this

OpenStudy (mathteacher1729):

Not that 1/2 = 3/6 and 1/3 = 2/6 The key is finding the common denominator 6. If you know how to find the least common multiple, then it's easy. 3 * 2 = 6 So the least common multiple is 6. for 2,4,8, and 5 The least common multiple is NOT 2*4*8*5... It is simply 8*5 = 40 That's because 8 is already divisible by 2 and 4.

OpenStudy (anonymous):

ok so the greatest common multiple is 40

OpenStudy (anonymous):

so then what do id with 40

OpenStudy (anonymous):

is this the common denominator of all of them

OpenStudy (mathteacher1729):

Yes. 40 needs to be the common denominator.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then i add all the numerators

OpenStudy (mathteacher1729):

No. You're looking for the middle number. When you have each fraction as (something) / 40 , line them up from smallest to biggest. The one in the middle is your answer. (you may have to simplify).

OpenStudy (anonymous):

ok so the numerators stay the same

OpenStudy (mathteacher1729):

No, they will change. Just as 1/2 changed to 3/6 both numerator and denominator of each fraction will change.

OpenStudy (anonymous):

ok

OpenStudy (mathteacher1729):

So 1 1/4 = 5 / 4 How do you write 5/4 as a fraction with denominator 40?

OpenStudy (anonymous):

8/40

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

15/40

OpenStudy (anonymous):

or 50/40

OpenStudy (anonymous):

right

OpenStudy (anonymous):

@mathteacher1729

OpenStudy (anonymous):

i get 20/40s then

OpenStudy (amistre64):

thats alot of work for something so simple

OpenStudy (amistre64):

sort the numbers from low to high

OpenStudy (amistre64):

1 1/4 , 5/8 , 3/5 , 1/2 , 1 1/2, 1 3/4 1/2, 3/5 , 5/8 , 1 1/4 , 1 1/2, 1 3/4 ^ this is the middle take the average of the middle "2" data points to get there

OpenStudy (anonymous):

ok thanks